# Sparse Bayesian Multidimensional Item Response Theory

### v3

@@ 1 Introduction
By explicitly characterizing latent factor posterior distributions, this paper develops a scalable Bayesian framework for estimating sparse exploratory MIRT models, with the aim of recovering an interpretable representation of the factor loading matrix $B$ without making assumptions on its zero allocations. The process of finding a sparse solution can be facilitated by the idea of parameter expansion [px-em], which enables rapid exploration of the parameter space of equivalent observed data likelihoods and is more robust against poor initialization. Our proposed framework and estimation procedures can also be further generalized beyond binary datasets to accommodate mixed continuous and ordinal responses. Parallel research in the CDM framework has also received significant attention recently: In CDMs, the latent structure, represented by the Q-matrix, is typically provided by experts. However, the Q-matrix can be misspecified or challenging to construct in practice. Consequently, sparsity recovery through the estimation of the Q-matrix has been explored in many recent works [slcm, qmat_dina, de_la_Torre2016].
@@
Besides its computational efficiency and versatility to accommodate mixed data type, one primary advantage of our Bayesian approach is its flexibility in prior specification. In many exploratory MIRT settings, the number of latent factors and the exact sparsity level of the loading matrix are typically unknown or can be misspecified by experts. To incorporate our prior beliefs amid such uncertainty, we impose a hierarchical Indian Buffet Process (IBP) prior [ibp] coupled with the Spike-and-Slab LASSO (SSL) prior [SSL] on the factor loading matrix, following the idea in [bfa]. This allows the algorithm to automatically learn the latent dimensionality and produce sparse loading estimates. Additionally, we provide general identification conditions under our proposed probit factor modeling framework and theoretical guarantees for recovering the true number of latent factors. When the number of observations is relatively small, which is often the case in social science and behavior health research, our Bayesian approach can yield more reliable estimates by leveraging such prior information, resulting in a more sparse solution.
@@
The rest of this paper is organized as follows. Section 2 offers a concrete motivating example exhibiting the potential of our proposed approach to estimating sparse Bayesian MIRT models. Section 3 provides a detailed exposition of the data generation process, identification conditions, and the theoretical justifications behind our proposed hierarchical model. We then show the derivations of our parameter-expanded EM algorithm in Section 4. Section 5 further demonstrates the generality of our proposed methodology by accommodating mixed data types, including continuous and ordinal responses. In Section 6, the effectiveness of our proposed method is substantiated through its application to diverse scenarios: a challenging synthetic dataset, a large sample educational assessment dataset, a quality-of-life measurement dataset, and a challenging ordinal bio-behavioral dataset. We conclude the paper in Section 7 and provide additional details in the Appendix. The relevant code with tutorials can be found in the following Github repo: https://github.com/JiguangLi/sparse_bayesian_mirt.
@@ 2 Exact Sparsity Recovery
For benchmarking, we consider the other three state-of-the art exploratory MIRT methods, where no assumptions are made for the zero allocations of the loading matrix : (1) the MHRM algorithm [mhrm_exploratory], (2) a customized Gibbs sampler with an adaptive spike-and-slab normal prior [ir_ss] (derivations available in Appendix A), and (3) the exploratory JML estimation method [jml1]. The MHRM and the JML method implementations are available in the mirt and the mirtjml R packages [JSSv048i06].
@@ 3 Nonparametric Bayesian MIRT
Building upon the Bayesian factor analysis framework introduced by Ročková and George [bfa], we deploy the spike-and-slab LASSO (SSL) prior [SSL] on the loading matrix to prompt sparsity. Sparsity plays an important role in our proposed model for multiple reasons: it alleviates model identification issues by anchoring estimation on sparse solutions, leading to more interpretable outcomes. Additionally, our application of the SSL prior reflects many real-world applications where factor loadings are inherently sparse, and therefore has the potential to improve out-of-sample prediction. Our SSL approach employs a two-point mixture of Laplace distributions which can adaptively produce exact zeros, and does not necessitate an additional thresholding step. For each element $B_{jk}$ in the loading matrix, we couple the SSL prior with a binary variable selection variable $\gamma_{jk}\in\{0,1\}$. Let $\lambda_{0k}>>\lambda_{1}>0$, and let $\psi(\beta|\lambda)=\frac{\lambda}{2}\exp\{-\lambda|\beta|\}$ denote the Laplace prior with mean 00 and variance $\frac{2}{\lambda^{2}}$. We consider the following hierarchical prior:
@@ 3.2 Posterior Asymptotics
The corresponding proof can be found in Section B.2. In the linear factor model setting [bfa], one may easily integrate Gaussian factor over the likelihood to study the posterior distribution of the covariance matrix. This approach is no longer possible under the probit link. Consequently, we introduce assumption (C) to control the Kullback-Leibler divergence and its variation between the posterior density and the true density.
@@ 4.2.1 Characterize Posterior Distribution of Latent Factors
The proofs of Theorem 4.2 and Corollary 4.3 can be found in Section C of the Appendix.
@@ 4.4 The PXL-EM Algorithm
We present the parameter-expanded version of our EM algorithm (PXL-EM) to expedite convergence and enhance exploration of the parameter space. The updates for our PXL-EM algorithm are almost as identical to vanilla EM algorithm described above, except we now need an additional auxiliary parameter $A$, which serves to rotate the loading matrix $B$ towards a sparser configuration (PXL-EM). A brief summary of our PXL-EM algorithm can be found in the table of Algorithm 1. The derivation and the justification of the factor rotation step is provided in the Appendix D.
@@ 5 Extension to Mixed Data Types
Theorem 5.1 is particularly valuable because it demonstrates that sampling from the latent factor posterior distribution remains tractable under mixed item response data types, as it still conforms to an instance of the unified skew-normal distribution. This ensures that our proposed E-step retains its efficiency and that our methodology can be generalized to handle a mix of continuous, binary, and ordinal item response types. The proof of Theorem 5.1, and the comprehensive EM procedure are provided in Section I of the Appendix.
@@ 6.2 IBP Loading Matrix
To assess whether PXL-EM can recover the true number of factors, we deliberately overshot the number of factors as $K^{*}=10$. Keeping $\lambda_{1}=0.5$ fixed, we gradually increased $\lambda_{0}$ along the path $\{0.5,1,3,6,10,20,30,40\}$ following the procedure described in Section 6.1. A visualization of the loading matrix estimation along this $\lambda_{0}$-path can be found in the Appendix E.1. Convergence was determined when $|B^{(t+1)}-B^{(t)}|_{\infty}<0.06$. Table 1 documents how the estimation accuracy dramatically improved over time.
@@
As shown in Table 2, our PXL-EM algorithm achieved the best overall performance in the accuracy of loading estimation, as well as the recovery of true zeros. One might have noticed that PXL-EM was not dramatically fast in this experiment. It is essential to realize our PXL-EM had already attained matching performance when $\lambda_{0}=6$. The Gibbs sampler would not scale well with either the sample size or the truncated dimension. The availability of more CPUs can always effectively accelerate our EM algorithm, but may not help speeding up the Gibbs sampler. As the sample size $N$ increases, we can expect the estimation of the factor loading matrix $B$ to become increasingly accurate. This is due to the IBP loading structure possessing a GLT structure, making the probit IBP factor model identifiable by Theorem 3.3. In Appendix E.2, we empirically verify consistency by repeating the same experiment with varying $N$ from $100$ to $1,000$.
@@ 6.4 Quality of Life Measurement
We further demonstrate the potential of our approach by applying it to a more intricate Quality of Life (QOL) dataset discussed in Lehman [Lehman1988AQO]. This dataset comprises responses from $586$ individuals with chronic mental illnesses, who answered $35$ items related to their life satisfaction across seven distinct subdomains (family, finance, health, leisure, living, safety, and social). In the psychological measurement literature, Gibbons et al. [bifactor_graded] showcased the advantages of the bifactor model in a confirmatory analysis. In their study, items associated with each subdomain loaded on separate secondary latent factors as well as the primary quality-of-life factor. We replicated Gibbons’ bifactor analysis, and the estimated bifactor loading structure is visualized in the leftmost plot of figure 5.
@@ 6.5 Ordinal bio-behavioral Data
In addition to identifying the negative correlation, we are also able to pinpoint the exact symptoms that are highly correlated with reduced cortical thickness. The research in [Stan2020] highlights seven symptom items with highly negative factor loadings lower than -0.25: delusions (PANSS-p1), hallucinatory behavior (PANSS-p3), suspiciousness (PANSS-p6), depression (PANSS-g6), unusual thought content (PANSS-g9), active social avoidance (PANSS-g16), and passive-apathetic social withdrawal (PANSS-n4). Remarkably, by examining the second bio-behavioral dimension, we identify six symptoms with factor loadings lower than -0.25, each of which is among the seven highlighted symptoms listed above. The only missing item is PANSS-g6, which still has a very negative loading of -0.24. We invite readers to refer to Tables 10 and 11 in the Appendix for the complete factor loading matrix and its comparison with the findings in [Stan2020].
@@ 7 Discussion
This article advocates for a novel Bayesian framework for estimating exploratory sparse MIRT models. The combination of the Indian Buffet Process prior and the spike-and-slab LASSO prior allows our EM algorithm to simultaneously learn latent dimensionality and sparse representations of factor loadings, which have not been sufficiently addressed in the binary and ordinal factor analysis literature. The efficiency of the algorithm is driven by the adaptation of parameter expansion, as well as by an efficient sampling strategy during the E-step, where we explicitly derived latent factor posterior distributions. We further demonstrate the advantage of adopting a probit link for establishing identification results and provide theoretical guarantees for recovering the true number of factors and contraction towards the true factor loading matrix under our prior specification. Our approach is capable of accommodating mixed data types, and we showcase the effectiveness of our approach on a challenging synthetic dataset and three real-world datasets from different domains. In each experiment presented in Section 6, our algorithm consistently generates sparse and interpretable estimations of the factor loading matrices, demonstrating its unique strength compared to other popular exploratory MIRT estimation frameworks.
While we focus on the estimation of the factor loadings and intercepts, our framework also has direct implications for latent trait measurements. For a well-calibrated set of $J$ items where the item parameters are known, Theorems 4.2 and 5.1 essentially provide an analytic form of the latent factor posterior distribution of any new observation $y\in\mathbbm{R}^{J}$. This avoids the additional computational need to perform MCMC sampling in order to conduct posterior inference on latent traits. Compared to the other MIRT estimation strategies, this can potentially save tremendous amount of computational time when the number of observations is large. Furthermore, many important properties of the unified skew-normal distribution such as the mean and covariance are readily available in [sun], making exact Bayesian inference possible. For more complicated objects of inferential interest, one may also consider the efficient posterior sampling strategies outlined in Section 4.2 to compute credible intervals. In Appendix G.2, we conducted a comparative analysis of the estimated latent factors and their posterior variances between our proposed PXL-EM approach and the oracle bifactor model for the QOL dataset, highlighting the potential of our approach in posterior factor inference and psychological measurements.
@@ Appendix B Sparse Bayesian MIRT: Model Properties
Part(b): Since the intercept is uniquely identified, we consider two sets of model parameters $(A,D)$ and $(B,D)$ defining the same probit factor model. We first observe if $\tilde{A}=\tilde{B}$, it must be the case $A=B$. This is because for any two K-dimensional vectors $a$ and $b$, the equation $\frac{a}{\sqrt{\|a\|^{2}+1}}=\frac{b}{\sqrt{\|b\|^{2}+1}}$ implies $a=b$. Hence it is sufficient to prove the theorem by showing $\tilde{A}=\tilde{B}$. Since $D_{1,j}=D_{2,j}$, equation 2 of proposition 3.1 implies $\|A_{j}\|=\|B_{j}\|$, which further implies $\Sigma_{A}=\Sigma_{B}$ as defined in equation 7. It follows we must have $\tilde{A}{A^{\prime}}=\tilde{B}{B^{\prime}}$. Since $A$ and $B$ conforms to an ordered GLT structure, we know $\tilde{A}$ and $\tilde{B}$ also follows an ordered GLT structure. Then we can simply apply Theorem 1 of [Fr_hwirth_Schnatter_2023] to obtain $\tilde{A}=\tilde{B}$, which implies $A=B$. ∎
@@ D.2 PXL-EM Updates
The E-step of our PXL-EM algorithm is the same as in section 4.2, except that we use $B=B^{*}A_{L}$ instead of $B^{*}$ when computing conditional expectation. For the M-step, the updates for $(B^{*},C,D)$ are the same as illustrated in Section 4.3. To update $A$, Consider the first-order condition:
@@ I.3.1 E-Step
Theorem 5.1 explicitly characterizes the latent factor posterior distribution under mixed data distributions as an instance of the unified-skewed normal distribution. The following corollary makes explicit the E-step.
@@
When the number of ordinal items is too large, it is possible to render the matrix $(\bar{D}\Omega_{\text{post}}\bar{D}^{\prime}+\bar{\mathbbm{I}}_{\bar{J}})$ non invertible, as both $\bar{D}\Omega_{\text{post}}\bar{D}^{\prime}$ and $\bar{\mathbbm{I}}_{\bar{J}})$ can become semi positive definite matrices. If this was the case, we may have to do a pseudo inverse. Alternatively, it is easy to see adding any $\epsilon>0$ on its diagonal would yield an invertible matrix. When we are just mixing continuous and binary items, the matrix $(\bar{D}\Omega_{\text{post}})\bar{D}^{\prime}+\bar{\mathbbm{I}}_{\bar{J}})$ is guaranteed to be positive definite, and hence invertible.

### v4

@@ 1 Introduction
By explicitly characterizing latent factor posterior distributions, this paper develops a scalable Bayesian framework for estimating sparse exploratory MIRT models, with the aim of recovering an interpretable representation of the factor loading matrix $B$ without making assumptions on its zero allocations. The process of finding a sparse solution can be facilitated by the idea of parameter expansion [px-em], which enables rapid exploration of the parameter space of equivalent observed data likelihoods and is more robust against poor initialization. Our proposed framework and estimation procedures can also be further generalized beyond binary datasets to accommodate mixed responses. Parallel research in the CDM framework has also received significant attention recently: In CDMs, the latent structure, represented by the Q-matrix, is typically provided by experts. However, the Q-matrix can often be misspecified in practice. Consequently, sparsity recovery through the estimation of the Q-matrix has been explored in many recent works [slcm, qmat_dina, de_la_Torre2016].
@@
Besides its computational efficiency and versatility to accommodate mixed data type, one primary advantage of our Bayesian approach is its flexibility in prior specification. In many exploratory MIRT settings, the number of latent factors and the exact sparsity level of the loading matrix are typically unknown and are often misspecified. To incorporate our prior beliefs amid such uncertainty, we impose a hierarchical Indian Buffet Process (IBP) prior [ibp] coupled with the Spike-and-Slab LASSO (SSL) prior [SSL] on the factor loading matrix, following the idea in [bfa]. This allows the algorithm to automatically learn the latent dimensionality and produce sparse loading estimates. Additionally, we establish general identification conditions and provide theoretical guarantees for recovering the true number of latent factors and ensuring the consistency of the loading matrix. When the number of observations is relatively small, which is often the case in social science and behavior health research, our Bayesian approach can yield more reliable estimates by leveraging such prior information, resulting in a more sparse solution.
@@
The rest of this paper is organized as follows. Section 2 offers a concrete motivating example exhibiting the potential of our proposed approach to estimating sparse Bayesian MIRT models. Section 3 provides a detailed exposition of the data generation process, identification conditions, and the theoretical justifications behind our proposed hierarchical model. We then show the derivations of our parameter-expanded EM algorithm in Section 4. Section 5 further demonstrates the generality of our proposed methodology by accommodating mixed data types. In Section 6, the effectiveness of our proposed method is substantiated through its application to diverse scenarios: a challenging synthetic dataset, a large sample educational assessment dataset, a quality-of-life measurement dataset, and a challenging ordinal bio-behavioral dataset. We conclude the paper in Section 7 and provide additional details in the Appendix.
@@ 2 Exact Sparsity Recovery
For benchmarking, we consider the other three state-of-the art exploratory MIRT methods, where no assumptions are made for the zero allocations of the loading matrix : (1) the MHRM algorithm [mhrm_exploratory], (2) a customized Gibbs sampler with an adaptive spike-and-slab normal prior [ir_ss] (derivations in Appendix A), and (3) the exploratory JML estimation method [jml1]. The MHRM and the JML method implementations are available in the mirt and the mirtjml R packages [JSSv048i06].
@@ 3 Nonparametric Bayesian MIRT
Building upon the Bayesian factor analysis framework introduced by Ročková and George [bfa], we deploy the spike-and-slab LASSO (SSL) prior [SSL] on the loading matrix to prompt sparsity. The sparsity assumption is essential, as it alleviates model identification issues by anchoring estimation on sparse solutions, leading to more interpretable outcomes. Additionally, the SSL prior reflects many real-world applications where factor loadings are inherently sparse, and therefore has the potential to improve out-of-sample prediction. Our SSL approach employs a two-point mixture of Laplace distributions which can adaptively produce exact zeros, and does not necessitate an additional thresholding step. For each element $B_{jk}$ in the loading matrix, we couple the SSL prior with a binary variable selection variable $\gamma_{jk}\in\{0,1\}$. Let $\lambda_{0k}>>\lambda_{1}>0$, and let $\psi(\beta|\lambda)=\frac{\lambda}{2}\exp\{-\lambda|\beta|\}$ denote the Laplace prior with mean 00 and variance $\frac{2}{\lambda^{2}}$. We consider the following hierarchical prior:
@@ 3.2 Posterior Asymptotics
The corresponding proof can be found in Appendix B.2. In the linear factor model setting [bfa], one may easily integrate Gaussian factor over the likelihood to study the posterior distribution of the covariance matrix. This approach is no longer possible under the probit link. Consequently, we introduce assumption (C) to control the Kullback-Leibler divergence and its variation between the posterior density and the true density.
@@ 4.2.1 Characterize Posterior Distribution of Latent Factors
The proofs of Theorem 4.2 and Corollary 4.3 can be found in Appendix C.
@@ 4.4 The PXL-EM Algorithm
We present the parameter-expanded version of our EM algorithm (PXL-EM) to expedite convergence and enhance exploration of the parameter space. The updates for our PXL-EM algorithm are almost as identical to vanilla EM algorithm described above, except we now need an additional auxiliary parameter $A$, which serves to rotate the loading matrix $B$ towards a sparser configuration (PXL-EM). A brief summary of our PXL-EM algorithm can be found in the table of Algorithm 1. The derivation and the justification of the factor rotation step is provided in Appendix D.
@@ 5 Extension to Mixed Data Types
Theorem 5.1 is particularly valuable because it demonstrates that sampling from the latent factor posterior distribution remains tractable under mixed item response data types, as it still conforms to an instance of the unified skew-normal distribution. This ensures that our methodology can be generalized to handle a mix of continuous, binary, and ordinal item response types. The proof of Theorem 5.1, and the comprehensive EM procedure are provided in Appendix I.
@@ 6.2 IBP Loading Matrix
To assess whether PXL-EM can recover the true number of factors, we deliberately overshot the number of factors as $K^{*}=10$. Keeping $\lambda_{1}=0.5$ fixed, we gradually increased $\lambda_{0}$ along the path $\{0.5,1,3,6,10,20,30,40\}$ following the procedure described in Section 6.1. A visualization of the loading matrix estimation along this $\lambda_{0}$-path can be found in Appendix E.1. Convergence was determined when $|B^{(t+1)}-B^{(t)}|_{\infty}<0.06$. Table 1 documents how the estimation accuracy dramatically improved over time.
@@
As shown in Table 2, our PXL-EM algorithm achieved the best overall performance in the accuracy of loading estimation, as well as the recovery of true zeros. One might have noticed that PXL-EM was not dramatically fast in this experiment. It is essential to realize our PXL-EM had already attained matching performance when $\lambda_{0}=6$. The Gibbs sampler would not scale well with either the sample size or the truncated dimension. The availability of more CPUs can always effectively accelerate our EM algorithm, but may not speed up the Gibbs sampler. As the sample size $N$ increases, we can expect the estimation becomes increasingly accurate. This is due to the IBP loading structure possessing a GLT structure, making it identifiable by Theorem 3.3. In Appendix E.2 , we empirically verify consistency by repeating the same experiment with varying $N$ from $100$ to $1,000$.
@@ 6.4 Quality of Life Measurement
We demonstrate the potential of our approach by applying it to a more intricate Quality of Life (QOL) dataset discussed in Lehman [Lehman1988AQO]. This dataset comprises responses from $586$ individuals with chronic mental illnesses, who answered $35$ items related to their life satisfaction across seven distinct subdomains (family, finance, health, leisure, living, safety, and social). In the psychological measurement literature, Gibbons et al. [bifactor_graded] showcased the advantages of the bifactor model in a confirmatory analysis. We replicated their analysis, and the estimated bifactor loading structure is visualized in the leftmost plot of figure 5.
@@ 6.5 Ordinal bio-behavioral Data
In addition to identifying the negative correlation, we are also able to pinpoint the exact symptoms that are highly correlated with reduced cortical thickness. The research in [Stan2020] highlights seven symptom items with highly negative factor loadings lower than -0.25: delusions (PANSS-p1), hallucinatory behavior (PANSS-p3), suspiciousness (PANSS-p6), depression (PANSS-g6), unusual thought content (PANSS-g9), active social avoidance (PANSS-g16), and passive-apathetic social withdrawal (PANSS-n4). Remarkably, by examining the second bio-behavioral dimension, we identify six symptoms with factor loadings lower than -0.25, each of which is among the seven highlighted symptoms listed above. The only missing item is PANSS-g6, which still has a very negative loading of -0.24. We invite readers to refer to the tables in Appendix H for the detailed estimates and comparison.
@@ 7 Discussion
This article advocates for a novel Bayesian framework for estimating exploratory sparse MIRT models. The combination of the Indian Buffet Process prior and the spike-and-slab LASSO prior allows our EM algorithm to simultaneously learn latent dimensionality and sparse representations of factor loadings, which have not been sufficiently addressed in the categorical actor analysis literature. The efficiency of the algorithm is driven by the adaptation of parameter expansion, as well as by an efficient sampling strategy during the E-step, where we explicitly derived latent factor posterior distributions. We further demonstrate the advantage of adopting a probit link for establishing identification results and provide theoretical guarantees for recovering the true number of factors and contraction towards the true factor loading matrix under our prior specification. Our approach is capable of accommodating mixed data types, and we showcase the effectiveness of our approach on a challenging synthetic dataset and three real-world datasets from different domains. In each experiment presented in Section 6, our algorithm consistently generates sparse and interpretable estimations of the factor loading matrices, demonstrating its unique strength compared to other popular exploratory MIRT estimation frameworks.
While we focus on the estimation of the factor loadings and intercepts, our framework also has direct implications for latent trait measurements. For a well-calibrated set of $J$ items where the item parameters are known, Theorems 4.2 essentially provides an analytic form of the latent factor posterior distribution of any new observation $y\in\mathbbm{R}^{J}$. This avoids the additional computational need to perform MCMC sampling in order to conduct posterior inference on latent traits. Compared to the other MIRT estimation strategies, this can potentially save tremendous amount of computational time when the number of observations is large. Furthermore, many important properties of the unified skew-normal distribution such as the mean and covariance are readily available in [sun], making exact Bayesian inference possible. For more complicated objects of inferential interest, one may also consider the efficient posterior sampling strategies outlined in Section 4.2 to compute credible intervals. In Appendix G.2, we conducted a comparative analysis of the estimated latent factors and their posterior variances between our proposed PXL-EM approach and the oracle bifactor model for the QOL dataset, highlighting the potential of our approach in posterior factor inference and psychological measurements.
@@
Acknowledgments
The authors sincerely thank the Associate Editor and four anonymous referees for their insightful comments and valuable suggestions, which have significantly enhanced the quality of this paper. Additionally, Robert D. Gibbons acknowledges support from the National institute on Aging (NIA-R56 AG084070), and Veronika Ročková acknowledges support from the National Science Foundation (NSF-DMS 1944740).
Disclosure Statement
The authors report there are no competing interests to declare.
@@ Appendix B Sparse Bayesian MIRT: Model Properties
Part(b): Since the intercept is uniquely identified, we consider two sets of model parameters $(A,D)$ and $(B,D)$ defining the same probit factor model. We first observe if $\tilde{A}=\tilde{B}$, it must be the case $A=B$. This is because for any two K-dimensional vectors $a$ and $b$, the equation $\frac{a}{\sqrt{\|a\|^{2}+1}}=\frac{b}{\sqrt{\|b\|^{2}+1}}$ implies $a=b$. Hence it is sufficient to prove the theorem by showing $\tilde{A}=\tilde{B}$. Since $D_{1,j}=D_{2,j}$, the equation in the Proposition 3.1 implies $\|A_{j}\|=\|B_{j}\|$, which further implies $\Sigma_{A}=\Sigma_{B}$ as defined in equation 7. It follows we must have $\tilde{A}{A^{\prime}}=\tilde{B}{B^{\prime}}$. Since $A$ and $B$ conforms to an ordered GLT structure, we know $\tilde{A}$ and $\tilde{B}$ also follows an ordered GLT structure. Then we can simply apply Theorem 1 of [Fr_hwirth_Schnatter_2023] to obtain $\tilde{A}=\tilde{B}$, which implies $A=B$. ∎
@@ D.2 PXL-EM Updates
The E-step of our PXL-EM algorithm is the same as in section 4.2.2, except that we use $B=B^{*}A_{L}$ instead of $B^{*}$ when computing conditional expectation. For the M-step, the updates for $(B^{*},C,D)$ are the same as illustrated in Section 4.3. To update $A$, Consider the first-order condition:
@@ I.3.1 E-Step
Theorem 5.1 explicitly characterizes the latent factor posterior distribution under mixed data distributions as a similar form to the unified-skewed normal distribution. The following corollary makes the E-step more explicit.
@@
When the number of ordinal items is large, it is possible to render the matrix $(\bar{D}\Omega_{\text{post}}\bar{D}^{\prime}+\bar{\mathbbm{I}}_{\bar{J}})$ non invertible, as both $\bar{D}\Omega_{\text{post}}\bar{D}^{\prime}$ and $\bar{\mathbbm{I}}_{\bar{J}}$ can become semi positive definite matrices. If this was the case, we may have to do a pseudo inverse. Alternatively, it is easy to see adding any $\epsilon>0$ on its diagonal would yield an invertible matrix, which can serve for a good approximation. In the bio-behavioral ordinal experiment, we added a small $\epsilon$ to the diagonal to ensure invertibility. When we are just mixing continuous and binary items, the matrix $(\bar{D}\Omega_{\text{post}}\bar{D}^{\prime}+\bar{\mathbbm{I}}_{\bar{J}})$ is guaranteed to be positive definite, and hence invertible.